A process takes 0.3 moles of helium gas from standard atmospheric pressure to 1.25 time its initial pressure. If we assume the process is isochoric and that the initial temperature is 16.5 Celsuis, what is the amount of heat supplied?
Answer #1

For isochoric process,

T2/T1 = P2/P1

T2 = 1.25* (273+16.5)

=361.87 K

Mass of helium,

m = n*mass number

=.3* 4 g

=1.2 grams

Heat supplied,

Q = m*Cv*(T2-T1)

= 1.2*10^-3*3.1156*10^3*(361.87-273-16.5)

=270.58 J.